It is the practice of almost all banks in the Philippines that when they grant a loan, the interest for one year is automatically deducted from the principal amount upon release of money to a borrower. Let us therefore assume that you applied for a loan with a bank and the P80,000 was approved at an interest rate of 14% of which P11,200 was deducted and you were given a check of P68,800. Since you have to pay the amount of P80,000 one year after, what then will be the effective interest rate?


1
Expert's answer
2022-03-17T14:11:26-0400

Let the principal (P)= 80000

Amount received (A)=68800

Interest rate (R) = 0.14

Total interest = 11200

Time for payment (T) =1 year

effective interest (i) = ?

total interest = "A*\frac{r}{100}*T"

11200 = 68800*"\frac{r}{100}*1"

1120000 =68800r

therefore r= "\frac{1120000}{68800}"

= 16.27906977

The effective interest rate is approximately 16.28%